OpenStudy (anonymous):

18. What is an equation for the line that passes through points (8, –3) and (9, 10)? (y + 3) = –13(x – 8) (y + 3) = 13(x – 8) (y – 3) = –13(x + 8) (y – 3) = 13(x + 8)

jimthompson5910 (jim_thompson5910):

What's the slope of the line that goes through (8, –3) and (9, 10)

OpenStudy (anonymous):

m = 13

OpenStudy (anonymous):

I think

jimthompson5910 (jim_thompson5910):

so m = 13

jimthompson5910 (jim_thompson5910):

oh you already said that, my bad

jimthompson5910 (jim_thompson5910):

because m = 13, you can plug this along with x1 = 8 and y1 = -3 into y - y1 = m(x - x1) to get y - y1 = m(x - x1) y - (-3) = 13(x - 8) y + 3 = 13(x - 8)

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw